### 17 Dec 2023 v3

[…]
${}^{1}$epartment of Statistics, USBE, Umeå University, Sweden.
${}^{2}$meå Center for Functional Brain Imaging, Umeå University, Sweden.
${}^{3}$epartment of Integrative Medical Biology, Umeå University, Sweden.
${}^{4}$epartment of Statistics, University of Florida, USA.
$*{}^{*}$orresponding author, E-mail: maria.josefsson@umu.se
Summary We investigate long-term cognitive effects of an intervention, where systolic blood pressure (sBP) is monitored at more optimal levels, in a large representative sample. A limitation with previous research on the potential risk reduction of such interventions is that they do not properly account for the reduction of mortality rates. Hence, one can only speculate whether the effect is a result from changes in cognition or changes in mortality. As such, we extend previous research by providing both an etiological and a prognostic effect estimate. To do this we propose a Bayesian semi-parametric estimation approach for an incremental intervention, using the extended G-formula. We also introduce a novel sparsity-inducing Dirichlet hyperprior for longitudinal data, demonstrate the usefulness of our approach in simulations, and compare the performance relative to other Bayesian decision tree ensemble approaches. In our study, there were no significant prognostic- or etiological effects across all ages, indicating that sBP interventions likely do not have a strong effect on memory neither at the population level nor at the individual level.
Key words: BART, dropout, G-computation, L-DART, missing not at random, deaths
1. Introduction
[…]
Two approaches for hypothetical interventions that only intervene on subjects that do not reaches a prespecified threshold are the threshold intervention [35] and the representative intervention (e.g. [27, 41]). Although both of these ensure treatment remains within a prespecified interval, it is common that a significant proportion of the population cannot control their systolic blood pressure (sBP) at optimal levels, due to, for example, poor, or absent, response to treatment (requiring the use of multiple medications). One alternative approach in these settings is a incremental intervention where a subjects treatment instead is shifted downward (or upwards) by some prespecified function [11, 25]. However, their approaches are not valid for longitudinal data with dropout and deaths.
Attrition due dropout and deaths is common in longitudinal aging studies, especially if subjects are followed over a long time period. The G-formula can address both missing response and missing covariate data under an assumption of missing at random (MAR; [32, 41, 18]). Strategies for sensitivity analyses have further been proposed in situations where the missingness is thought to be missing not at random (MNAR) [14, 15]. However, cognition after death is usually not defined and not of interest to study, prompting consideration of a mortal cohort approach [7]. Mortal cohort inferences truncate participants’ outcomes after death. The first approach, partly conditional inference (PCI), condition on the sub-population who are still alive at that time point [19, 38, 37], the second approach, principal stratification, condition on the subpopulation surviving irrespective of treatment [8, 9]. Although the two approaches have similarities, they differ with respect to what type of information they generate. PCI has been proposed for estimating non-causal associations and is particularly useful for prognostication, i.e. when interest is in the health outcome among all individuals who are alive. The approach has for example been considered when intervening on risk factors for coronary heart disease [35]. In contrast, principal stratification is useful in situations when survival differ between treated and controls, and has been applied when interest is in etiological effect estimates [12, 36, 34, 14]. Since the main focus of previous intervention studies was to make prognostic rather than etiological intervention estimates, they have only considered PCI (or simply immortal cohorts) and principal stratification has not been adapted to this setting.
In the present study we investigate long-term cognitive effects of an sBP intervention, using prospective cohort data from a large representative sample [26]. An important limitation in prior research is the lack of sufficient consideration of altered mortality rates in evaluating the potential risk reduction. As such, we propose both partly conditional inference and inference on the principal strata. Thus, our study extends previous research by providing both etiological and prognostic intervention effect estimates, which are both critical for public health practice. In particular, we will be able to answer the following research questions; Would the proposed sBP intervention affect cognition levels in the general population? and Would the proposed intervention affect cognition at the individual level? These will offer a more comprehensive understanding of the intervention’s impact. For estimation, we propose a Bayesian semi-parametric approach based on the extended G-formula [31]. In addition, we introduce a novel sparsity-inducing Dirichlet hyperprior designed for regression with longitudinal data, and demonstrate the usefulness of our approach in a simulation relative to other Bayesian decision tree ensemble approaches.
The paper is organized as follows. In Section 2, we present the Betula data. In Section 3, we introduce the proposed intervention in a Bayesian setting. In Section 4, we describe the two types of mortal cohort inferences. In Section 5, we provide computational details and introduce a sparsity-inducing Dirichlet hyperprior for longitudinal data. In Section 6, we conduct a simulation study to assess the operating characteristics of the new prior, and in Section 7, we apply the methodology to the Betula data to answer the substantive question of interest. We close in Section 8 with some conclusions.
2 A blood pressure intervention
The subjects in this study are all randomly recruited participants in the Betula prospective cohort study, a population-based study with the objective to study memory, aging and dementia [26]. We consider 15-years of longitudinal data and four waves of data collection, administered in approximately five years intervals, where the first wave was initiated in 1988-1990. A total of n=800 participants were included, 100 participants from each of the eight age-cohorts: $35,40,\ldots,70$ years old at enrollment.
The cognitive outcome in this study was assessed using a composite episodic memory task, with a range of 0 to 76 ([16]). We consider a set of baseline confounding factors; age at enrollment, sex, and years of education, a set of time-varying cardiovascular risk factors; Body mass index (BMI), smoking, diastolic blood pressure, serum-cholesterol, and self reported type 2 diabetes, as well as, previous episodic memory measures. See Table A3 for baseline summary statistics. 38% dropped out of the study before the 15 year follow-up and of these 18% died or developed dementia.
[…]
3 Methods
Consider a prospective cohort study where measurements on a set of variables are collected on each of $i=1,2,\ldots,n$ subjects over a specified follow-up period at $t=1,2,\ldots,T$ fixed time points. Let $L_{it}$ denote the vector of possibly time-varying confounders and $A_{it}$ denote a continuous treatment measured at time $t$. Denote the continuous outcome measure by $Y_{it}$, where interest in this study is the outcome measure at the last test wave $T$, i.e. $Y_{iT}$. Furthermore, let $S_{it}=1$ indicate if a subject is alive at time point $t$ and $S_{it}=0$ otherwise. Let $R_{it}=1$ indicate if subject $i$ has completed the testing at time $t$, otherwise $R_{it}=0$. We have monotone missingness, so if $R_{it}=0$, $R_{ik}=0$ for $k>t$. Furthermore, let $T^{r}_{i}$ denote the number of time points a subject participates in the study, and $T^{s}_{i}\geq T^{r}_{i}$ denotes the number of time points the subject is alive. Let the observed history of a time-varying variable be denoted with an overbar and the future with an underbar. We also write $H_{i}$ to denote all the observed past history just prior to the outcome variable at the last time point $T$, such that $H_{i}=(\bar{Y}_{it-1}=\bar{y}_{it-1},\bar{A}_{it}=\bar{a}_{it},\bar{L}_{it}=\bar{l}_{it},\bar{R}_{it}=\bar{r}_{it},\bar{S}_{it}=1)$, with support on the set of possible histories $\mathcal{H}_{i}$ and where lower-case letters denote possible realizations of a random variable. At each time point $t$, we assume data is observed in the temporal ordering $S_{it},R_{it},L_{it},A_{it},Y_{it}$, and for each subject we observe $(\bar{s}_{iT^{s}_{i}},\bar{r}_{iT^{r}_{i}},\bar{l}_{iT^{r}_{i}},\bar{a}_{iT^{r}_{i}},\bar{y}_{iT^{r}_{i}})$. Below, we suppress the subscript $i$ to simplify notation.
3.1 A hypothetical incremental intervention on the NVT
[…]
Here we propose an incremental intervention [11, 25] based on the NVT, where treatment is shifted at time $t$ based on a parameter $\delta_{t}$. In a Bayesian setting it is possible to quantify the uncertainty in the intervention by imposing a prior on the shift parameter $\delta_{t}$, $p^{\delta_{t}}(\cdot)$, reflecting the researchers beliefs about the possible measures for treatment to take under the hypothetical intervention. We describe possible priors in relation to the Betula data in Section 7.
Similar to a threshold intervention [35], we one only intervene on subjects whose NVT at time $t$ does not meet a particular threshold, here denoted by $\tau$. However, instead of assigning treatment to the exact threshold value, $\tau$, treatment is assigned using the prior, $p^{\delta_{t}}(\cdot)$. Specifically, if $a_{t}^{*}\leq\tau$, then $p^{int^{*}}(a_{t}|\bar{y}_{t-1},a_{t}^{*},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)=1,$ if $a_{t}=a_{t}^{*}$. However, if $a_{t}^{*}>\tau$, then
$p^{int^{*}}(a_{t}|\bar{y}_{t-1},a_{t}^{*},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)=p^{\delta_{t}}(\delta_{t})I\{a_{t}=a_{t}^{*}-\delta_{t}\}.$
4 Intervention effects for mortal cohorts
Below we describe two types of mortal cohort inferences under the proposed incremental intervention on the NVT: partly conditional inference and inference on the principal strata. For simplicity, we temporarily restrict our description to the regimes $g\in(g^{int^{*}},g^{nc})$, where $g^{int^{*}}$ is the set of all deterministic intervention regimes based on the NVT, with support on the set of possible regimes $\mathcal{G}$, and $g^{nc}$ denote the natural course, i.e. the natural progression as if no intervention was considered. Furthermore, let $\bar{Y}_{T}(g)$, $\bar{A}_{T}(g)$, $\bar{R}_{T}(g)$, $\bar{S}_{T}(g)$, and $\bar{L}_{T}(g)$ represent the counterfactual outcome, treatment, missingness, survival, and covariate histories associated with the regime $g$.
4.1 Partly conditional inference
Partly conditional models have been proposed to address truncation by death under hypothetical interventions. For these models inference is conditioning on the sub-population being alive at a specific time-point, in here, being alive at end of follow-up, i.e. time $T$. Hence, inferences focus on what we call a partly conditional intervention effect ($PCIE$),
$PCIE=E[Y_{T}(g^{int^{*}})|S_{T}(g^{int^{*}})=1]-E[Y_{T}(g^{nc})|S_{T}(g^{nc})=1].$ (1)
Although the $PCIE$ gives a crude comparison of differences in potential outcomes under the two regimes, which is useful for prognostication, it often compares potential outcomes for different underlying populations. For example, it is likely that an sBP intervention may also affect survival rates. As such, the population surviving under the natural course will differ from the one surviving under the intervention.
4.2 Assumptions and identification of the PCIE
We now define a set of identifying assumptions for the $PCIE$.
C1 Consistency: For $g\in(g^{int^{*}},g^{nc})$ and $t=1,\ldots,T$, if $\bar{A}_{t}=\bar{A}_{t}(g)$, then $\bar{Y}_{t}=\bar{Y}_{t}(g)$, $\bar{R}_{t}=\bar{R}_{t}(g)$, $\bar{S}_{t}=\bar{S}_{t}(g)$ and $\bar{L}_{t}=\bar{L}_{t}(g)$.
C2a Conditional exchangeability under the natural course: For $g^{nc}$ and $t=1,\ldots,T$, conditional on the history, the treatment and the potential outcome are independent. That is $\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and $\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$.
C2b Conditional exchangeability for NVT interventions: For $g^{int^{*}}$
$\left(\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muA\mkern-3.0mu}\mkern 3.0mu_{t}^{*}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and $\left(\mkern 3.0mu\underline{\mkern-3.0muL\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$. The assumption states that under the intervention, the NVT at time $t$ has no direct effect on the cognitive outcome, or survival, except through future treatment measures. Hence, an implication of this assumption is that the NVT is not itself a time-varying confounder [29, 40].
C3 Sequential positivity: If $p(\bar{y}_{t-1},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)\neq 0$ then $p(a_{t}|\bar{y}_{t-1},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)>0$ for all $\bar{a}_{t}$ consistent with regime $g^{int^{*}}$.
C4 Missingness at follow-up: For the outcome, treatment and time-varying confounders we assume missing at random conditional on survival (MARS) on the distribution. In particular, for all $t>1$, $p(y_{t}|\bar{a}_{t},\bar{l}_{t},r_{t}=0,\bar{s}_{t}=1,\bar{y}_{t-1})=p(y_{t}|\bar{a}_{t},\bar{l}_{t},\bar{r}_{t}=1,\bar{s}_{t}=1,\bar{y}_{t-1})$, $p(a_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=0,\bar{s}_{t}=1,l_{t})=p(a_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=1,\bar{s}_{t}=1,l_{t})$, and $p(l_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=0,\bar{s}_{t}=1,l_{t-1})=p(l_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=1,\bar{s}_{t}=1,l_{t-1})$.
For the Betula data we perform a set of sensitivity analyses where we investigate robustness of results to violations of assumptions C2, C3, and C4. See Section 7.2.
For longitudinal cohort data, under the assumptions C1, C2a, C3 and C4, the mean outcome at time $T$ under the natural course, $g^{nc}$, can be computed using the g-formula (Robins, 1986) conditioning on survival. That is
$\int_{\bar{s}_{T}}\int_{\bar{r}_{T}}\int_{\bar{l}_{T}}\int_{\bar{a}_{T}}\int_{\bar{y}_{T-1}}E[Y_{T}|\bar{y}_{T-1},\bar{a}_{T},\bar{l}_{T},\bar{r}_{T},\bar{s}_{T}=1]$
×$\quad\times\prod_{k=1}^{T}p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$
×$\quad\times p(l_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k},\bar{s}_{k}=1)$
×$\quad\times p(r_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k}=1)$
×$\quad\times p(s_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1)$
×$\quad\times p(y_{k-1}|\bar{y}_{k-2},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1).$ (2)
To compute (S0.Ex2), Monte Carlo integration is implemented over the set of temporally preceding variables under the natural course, denoted by $H_{nc}$, among those who are alive at the final time point $\bar{s}_{T}=1$. Note, $s_{T}=1$ must also be sampled for.
In contrast, for the intervention, $g^{int^{*}}$, under assumptions C1, C2b, C3 and C4, the mean outcome at time $T$ for $g^{int^{*}}$ can be computed using the Extended g-formula [31]. The main difference is that $p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$ in (S0.Ex2) is replaced by $\int_{a_{t}^{*}}p^{int^{*}}(a_{k}|\bar{y}_{k-1},a^{*}_{k},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)\times p(a^{*}_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)$; details are given in Supplementary Appendix A. Similar to (S0.Ex2), Monte Carlo integration is implemented over the history under the intervention, denoted by $H_{int^{*}}$, among those who are alive. However, instead of sampling pseudo treatment data from the observed data distribution, as in (S0.Ex2), pseudo NVT data is first sampled from the observed data distribution. For those whose NVT does meet the threshold $\tau$, treatment is assigned to the NVT, and among those whose NVT does not meet the threshold $\tau$, treatment is assigned such that the NVT is shifted downwards by $\delta_{t}$, as described in Section 3.2.
The $PCIE$ in (1) is then computed by subtracting the expected outcome under regime $g^{nc}$ from the expected outcome under regime $g^{int^{*}}$. More details on identification are found in Supplementary Appendix B.
4.3 Hypothetical interventions on the principal strata
Conditioning on survival as described in Section 4.1 may introduce bias due to the fact that survival is a post-randomization event [42]. Hence, the partly conditional approach described cannot properly address etiological research questions in mortal cohorts. One alternative estimand is the survivors average causal effect, i.e. the causal effect on the subpopulation surviving under both regimes, $g^{nc}$ or $g^{int^{*}}$, i.e. the principal strata ([8, 9]). Our goal here is to estimate the difference in expected potential outcomes between $g^{nc}$ and $g^{int^{*}}$ on the principal strata. That is, a survivors average intervention effect ($SAIE$),
$SAIE=E[Y(g^{int^{*}})-Y(g^{nc})\mid S_{T}(g^{nc})=S_{T}(g^{int^{*}})=1].$ (3)
Note, previous studies have shown a the link between reduced sBP and reduced all-cause mortality [2]. Hence, it is likely that an intervention targeting sBP not only may have an effect on cognition but also on mortality rates. In such situations, the $PCIE$ would compare potential outcomes for different underlying populations. In particular, if subjects live longer under the risk-reducing intervention this population may include less healthy subjects who would die under the natural course, masking a possible intervention effect. The $SAIE$ on the other hand, compare potential outcomes from the same population, i.e. a population consisting of subjects who would survive both when intervening on sBP and under the natural course. Thereby, providing a estimate that adjust for differences in mortality rates for the two regimes.
In our application we are interested in both a prognostic effect estimate (i.e. the $PCIE$) and an etiological effect estimate (i.e. the $SAIE$). For identification of the $SAIE$, we need two additional assumptions.
C5 Stochastic Monotonicity: $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]\geq\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=0]$. Given a beneficial intervention, we assume the probability of survival under the intervention $g^{int^{*}}$ is higher among those subjects that would survive under the natural course $g^{nc}$, compared to those who would not. We further let $\Pr[S_{T}(g)=1]=\psi^{g}$, for $g=(g^{nc},g^{int^{*}})$, and assume $\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]=\psi^{g^{int^{*}}}+\lambda\left[\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}-\psi^{g^{int^{*}}}\right]$, where $\lambda$ is a sensitivity parameter and $0\leq\lambda\leq 1$ [33, 21]. Note, if $\lambda=1$ this assumption corresponds to a deterministic monotonicity assumption where subjects who were to be alive under the observed regime $g^{nc}$ would also be alive under the intervention $g^{int^{*}}$, i.e. $S_{T}(g^{nc})\leq S_{T}(g^{int^{*}})$. In contrast, if $\lambda=0$ this corresponds to an assumption of $S_{T}(g^{nc})$ and $S_{T}(g^{int^{*}})$ being independent.
C6 Difference in expectations of outcomes when comparing different strata. For $g\in(g^{int^{*}},g^{nc})$, $g^{\prime}\in(g^{int^{*}},g^{nc})$ and $g\neq g^{\prime}$, we assume, $E[Y_{T}(g)|S_{T}(g^{\prime})=S_{T}(g)=1]-E[Y_{T}(g)|S_{T}(g)=1,S_{T}(g^{\prime})\neq 1]=\Delta^{g}$ and further that $\Delta^{g}=\Delta^{g^{\prime}}=\Delta$. Note, if $\Delta=0$ there is no difference in potential outcomes when comparing the ”always survivor” strata to the strata where individuals were to live under regime $g$ but not under the contrasting regime $g^{\prime}$. In contrast, if $\Delta>0$ this implies a higher potential outcome is observed for the principal strata, i.e. the strata where individuals were to live under both regimes.
A detailed description of identification of the $SAIE$ is given in Supplementary Appendix B. In short, under assumptions C1-C6, the $SAIE$ is given by:
$\quad PCIE+\Delta\left\{\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)\right\}\left(1-\frac{1}{U}\right),$ (4)
where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. In particular, under A5 (stochastic monotonicity) (4) represents a location shift upwards from the $PCIE$ for a beneficial intervention where we expect the participants to improve their health under the intervention regime compared to the natural course. Note, $\psi^{g}=E_{H^{g}}\left\{\Pr[S_{T}=1|H^{g}]\right\}$, for $g\in(g^{int^{*}},g^{nc})$, are obtained by marginalizing over the distributions of the set of temporally preceding variables $H^{nc}$ using the g-formula and $H^{int^{*}}$ using the extended g-formula.
5 A semi-parametric approach for estimation
[…]
As for traditional (hard) BART models, we have that the mean function of a variable conditional on the history, is given by the sum of regression trees, $\mu_{\cdot}=\sum_{t=1}^{T}g_{y_{t}}(\mathcal{T}_{t},\mathcal{M}_{t})$, where $\mathcal{M}_{t}$ is the parameters for the leaf nodes and $\mathcal{T}_{t}$ the tree structure. In general, the distribution of the continuous variable are moreover specified as normal, and the binary variables are specified as probit models.
The regularization prior, controlling the complexity of the sum-of-trees, differ in two ways between the standard BART algorithm and soft BART. First, for soft BART, the path down the decision tree is probabilistic, where the sharpness of the decision is controlled by the bandwidth parameter. Thereby adapting to the smoothness level of the regression function even for low dimensional tree ensembles. Second, a predictor $x_{j}$, is chosen according to a probability vector $\mathbf{q}=(q_{1},\ldots,q_{P})$, typically specified as uniform. In contrast, soft BART models incorporate a Dirichlet distribution for $\mathbf{q}$, such that $\mathbf{q}\sim\mathcal{D}\left(\alpha/P,\ldots,\alpha/P\right)$. The sparsity-inducing prior was first introduced for high dimensional settings [22], and can be extended to allow sparsity both within and between groups [23, 6]. Determining the degree of sparsity in the model is done by placing a prior on $\alpha$, as such, $\frac{\alpha}{\alpha+\rho}\sim Beta\left(a,b\right)$. Default specifications for $a,b$ and $\rho$ is $0.5$, $1$ and $P$ respectively. This sparsity inducing prior is implemented for both hard and soft BART. We refer to BART models with this prior as Dirichlet additive regression trees (DART).
5.2 A grouping prior for longitudinal data
In a longitudinal setting, one may want to introduce a larger degree of sparsity for predictors collected at earlier time points, if they are thought to be less strongly associated with the response. Below we introduce a novel sparsity-inducing prior.
We consider a hierarchical structure where the predictors are divided into two groups: predictors measured at the current time point $t$ and predictors measured at past time points $k=1,\ldots,t-1$. Moreover, baseline predictors, e.g. age and sex, are grouped with the current predictors into one group, since we want to impose the same degree of sparsity on these predictors. Let $w$ refer to the probability of a predictor being from the set of current predictors, versus from the set of past predictors; we specify a beta prior, $w\sim Beta\left(a,b\right)$. For the current predictors, let $v^{t}_{j}$ refer to the probability of selecting the $j$th predictor from the set of current predictors $\textbf{v}^{t}=(v^{t}_{1},\ldots,v^{t}_{P_{t}}$); $\textbf{v}^{t}$ is given a Dirichlet prior $\textbf{v}^{t}\sim\mathcal{D}\left(\eta/P_{t},\ldots,\eta/P_{t}\right)$. So a priori, the probability of selecting the $j$th current predictor to construct a given split is equal to $w\times v^{t}_{j}$. The past predictors are divided into $t-1$ groups, where $P_{k}$ is the size of group associated with time $k$. Let $u^{k}$ denote the probability of selecting the $k$th set of time-varying covariates, and let $v^{k}_{j}$ denote the probability of selecting the $j$th predictor in the $k$th set of time-varying covariates. To introduce sparsity related to proximity in time to the response we give u a grouping Dirichlet prior $\textbf{u}\sim\mathcal{D}\left(\alpha_{1}/t-1,\ldots,\alpha_{t-1}/t-1\right)$, where we determine the degree of sparsity in the model by placing a prior on $\alpha_{k}$ such that a larger degree of sparsity is obtained for smaller $k$’s (see details below). Similarly $\textbf{v}^{k}$ is also given Dirichlet prior $\textbf{v}^{k}\sim\mathcal{D}\left(\varphi^{k}/P_{k},\ldots,\varphi^{k}/P_{k}\right)$. So a priori, the probability of selecting the $j$th predictor from the $k$th time point to construct a given split is equal to $(1-w)\times u^{k}\times v^{k}_{j}$.
For the hyperparameters $\eta$ and $\varphi^{k}_{j}$ we use the default Beta hyperprior as implemented in DART. For $\alpha_{j}$ we specify a Beta hyperprior $\frac{\alpha_{j}}{\alpha_{j}+\rho_{j}}\sim Beta\left(c_{j},d_{j}\right)$; we let $c_{j}$ vary as a linear function of time, set $d_{j}=1$, and $\rho_{j}=t$ for all $j$. To introduce sparsity related to proximity in time to the response we set $c_{j}=1-\frac{(t-j)}{t-1}*0.5$. This results in ordered hyperpriors for the $\alpha_{j}$s, where $c_{1}=0.5$ for predictors measured at the first time point and $10.5<c_{j-1}<c_{j}<$ 1 for $j>1$. That is, a higher degree of sparsity is introduced for predictors measured earlier in time and less sparsity for predictors measured closer in time to the response. The simulations in Section 6 shows the strong performance of these hyperprior choices. Details about the posterior computations of the L-Dirichlet prior are obtained in the Supplementary Appendix D.
5.3 Algorithm to compute the $PCIE$ and $SAIE$
The $PCIE$ and the $SAIE$ can be computed using the algorithm described in the Supplementary Appendix D and is implemented in the R package GcompLDART (available on https://github.com/m4ryjo/GcompLDART).
6 Simulation study
In this simulation study we evaluate the performance of L-DART and compare it to other BART specifications. Data were generated according to a modified version of the Friedman’s example [10], mirroring a longitudinal study with five time-varying predictors and a continuous outcome, with $f(\bar{\mathbf{x}}_{i})=\sum_{t=1}^{T}c_{t}[10\mathrm{sin}(\pi x_{t}^{1}x_{t}^{2})+20(x_{t}^{3}-0.5)^{2}+10x_{t}^{4}+5x_{t}^{5}]$, where $Y_{iT}\sim\mathrm{Normal}\left\{f(\bar{\mathbf{X}}_{i}),\sigma_{T}\right\}$, $\sigma_{T}=10$, and $c_{t}$ determines the strength of association between the predictors at time $t$ and the outcome.
The longitudinal predictors, $\bar{\mathbf{X}}_{i}=\bar{X}_{i}^{1},\ldots,\bar{X}_{i}^{T}$, were generated as correlated uniform random variables, generated as multivariate normal variables, with zero mean and covariance as described below, and then transformed to uniform variables. As such, $\bar{X}_{it}^{p}\sim Uniform(0,1)$, where the between time-points covariance for predictor $X^{p}$ is $Var(X^{p}_{it},X^{p}_{ik})=d_{tk}^{2}$. Here, we specified $d_{jj-1}=0.4$, $d_{jj-2}=0.2$ and $d_{jj-3}=0.1$, reflecting a stronger correlation for measures closer in time, and all other were set to zero.
We simulated 500 datasets of size $n=\left(100,500,1500\right)$, and five different scenarios by varying $\bar{c}=\left(c_{1},\ldots,c_{T}\right)$: Scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, reflecting a decreasing, although a relatively strong, association over time between the predictors and the outcome; Scenario B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, i.e. a sharp decrease in association over time; Scenario C: $\bar{c}=(1,0,0,0)$, i.e. only predictors at the closest time point are associated with the outcome; Scenario D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, which is similar to Scenario A but also includes four time points uncorrelated with the outcome; and Scenario E: $\bar{c}=(0.25,0.25,0.25,0.25)$, i.e. a weak and stable association with the outcome for all time points. For scenarios A, B, C, and, E, $T=4$ and for scenario D, $T=8$. For each scenario, 10 or 100 trees were specified for each model.
We compared the predictive performance of our proposed approach (L-DART) to two hard BART models; the standard BART algorithm (H-BART) and BART with the sparsity-inducing Dirichlet prior (H-DART) (as implemented in the R package BART), and two soft BART models; the standard S-DART algorithm (as implemented in the R package SoftBart) and SoftBart incorporating only the predictors from the last time point in the model (S-DART-lag1). For evaluation of their performance, the methods were first trained on one dataset and compared on a test dataset using: % bias, root-mean-squared error (RMSE), and 95% coverage probabilities (CPs), for all the predictions of the test data. Computed as the posterior mean of: %$\%bias=\frac{1}{\sum_{i=1}^{n}\mu_{T}}\sum_{i=1}^{n}\frac{y_{iT}-\mu_{iT}}{n}$, where $\mu_{T}=\frac{1}{n}\sum_{i=1}^{n}\mu_{iT}$ and $\mu_{iT}$ are obtained from the training data; $RMSE=\sqrt{\sum_{i=1}^{n}\frac{(y_{iT}-\mu_{iT})^{2}}{n}}$; and the $95\%$ CP is the average pointwise predictive coverage of $95\%$ CIs for all the predictions of the test data set. The models use their respective default priors (if not stated otherwise) and were fitted by using 3000 warm-up iterations, 400000 sampling iterations, and a thinning of 100. All simulations were done using R-4.1 at the High Performance Computing Center North (HPC2N) and code is available on https://github.com/m4ryjo/GcompLDART.
6.1 Simulation Results
All models were generally unbiased or nearly unbiased across all scenarios, and as expected, the biases/RMSEs were smaller and the coverage probabilities (CPs) were closer to their true values for larger sample sizes; see Table 1 and Table A1 for details.
The soft BART models using the full set of predictors, had generally the smallest RMSEs, and CPs close to the nominal values, with an advantage for L-DART. However, under scenario B and C (a strong association between predictors at the last time point while the association for the history is weak or zero), L-DART had much smaller RMSEs compared to S-DART. Moreover, in B and C for the larger sample sizes, as expected, the soft BART model using only the last time-point predictors (S-DART-lag1) performed best in terms of RMSEs and CPs, but by small margin over L-DART. For the smaller sample size (n=100), all CPs $\leq 0.91$ for S-DART-lag1. For all other scenarios, S-DART-lag1 produces RMSEs that are notably larger than the other models and with CPs well below the nominal level, $0.72-0.86$.
Moreover, for all scenarios, the RMSEs for the hard BART models were smaller when fitting 100 regression trees compared to 10, while producing overly conservative credible intervals. In contrst, when fitting 10 trees, CPs were closer to their true values. No improvement in RMSEs and CPs were generally seen for the soft BART models by increasing the number of trees, and the results did not differ significantly from fitting soft BART models with only one tree (Table A2).
7 Analysis of the Betula data
We consider an intervention where we assign sBP by following the incremental threshold intervention described in Sections 2 and 3. Here, if a subjects’s NVT at times $t=2,\ldots,T$ is greater than $\tau=140$, then sBP is assigned by imposing a incremental shift on the NVT, and otherwise we do not intervene on this subjects. Moreover, we specify a triangular prior for the shift parameter $\delta_{t}$, such that $\delta_{t}\sim Triangular(L_{\tau},0,L_{\tau})$. The lower bound are given by; $L_{\tau}\approx 44$ mm Hg, corresponding to two standard deviations of baseline sBP.
To obtain an estimate of the $SAIE$, we incorporate a prior on the sensitivity parameters in Assumptions C5 and C6. We again specify triangular priors for the sensitivity parameters $\Delta$ and $\lambda$. We let $\Delta\sim\mathrm{Triangular}(0,U_{\Delta},U_{\Delta}),$ with upper bound (and mode) set to $4.6\%$. Here, $4.6\%$ corresponds to the estimated difference in baseline memory between those who are alive at time point $T$ and those who are dead, when adjusting for age and sex; it is unlikely the difference between the strata would to be bigger than this. We further let $\lambda\sim Triangular(0.5,1,1)$, where the upper bound corresponds to a deterministic Monotonicity assumption while the lower bound reflects a moderate correlation between survival under either regimes.
7.1 Main results
We estimated the $PCIE$ and the $SAIE$ using L-DART with priors on the sensitivity parameters as described in the previous section. For each chain we used a thinning of 100, the first 300 iterations were discarded as burn-in, and 2,400 posterior samples were kept. We sampled pseudo data of size $N^{*}=10,000$ at each iteration for the MC integration. Convergence of the posterior samples was monitored using trace plots. Note, all results are given in terms of % correct scores of the maximum (i.e. 76).
Results for the different age-cohorts are given in Figure 1. The results revealed that posterior mean episodic memory score decreased with advancing age both under the natural course and under the hypothetical intervention, although, to a lesser extent for the latter. Among subjects 70 and older, this shift corresponds to postponing memory decline approximately $1.5$ years on average, although the effect is non-significant. The results for the $PCIE$ suggests that there is no long-term cognitive effects of monitoring sBP in this population (credible intervals cover 0 for all age-cohorts), although, the effect is notably stronger for the older age-cohorts compared to middle-aged. The estimates of the $PCIE$ and the $SAIE$ were similar. For the $SAIE$ this suggests that there is no long-term cognitive effects among those surviving irrespective of intervention. Hence, the non-significant results for the $PCIE$ are not driven by altered mortality rates. We moreover note that only small differences were found when comparing survival probabilities under the natural course and under the hypothetical intervention for all age cohorts, and as a result, leading to rather small differences between the $PCIE$ and the $SAIE$.
7.2 Sensitivity analyses
Unmeasured confounding: We implement a simulation procedure to account for an unmeasured time-varying confounder missing for time-points $k>l$. In the Betula study, measures for cholesterol are obtained only for the first two time points; hence, it is a possible unmeasured confounder for $k>2$. In the first step, $U_{k}$, for $k>2$, is drawn from an estimate of the conditional distribution of $p(u_{k}\mid H^{g}_{u_{k}})$. Here, we assume $p(u_{k}\mid H^{g}_{u_{k}})=p(u_{2}\mid H^{g}_{u_{2}})$, where $U_{2}$ is the last observed time point for $U_{k}$ and $H^{g}_{u_{2}}=(\bar{l}_{2},\bar{r}_{2},a_{1},y_{1})$ is the history for $U_{2}$. To obtain a plausible fit for the conditional distribution of $p(u_{k}\mid H^{g}_{u_{k}})$, this involves first obtaining a fit for $p(u_{2}\mid H^{g}_{u_{2}})$ using L-DART, then plugging $H^{g}_{u_{k}}$ into the L-DART model (for $u_{2}$), and, finally, drawing realizations for $U_{k}$ from the corresponding conditional distribution. Once we obtain realizations of $U_{k}$ for $k>2$, we then fit the (extended) g-formula to the full data, and estimate the intervention effect as described in Section 4. Since $H^{g}_{u_{2}}$ and $H^{g}_{u_{k}}$ corresponds to different number of predictors, we incorporate only a subset of the predictors for $H^{g}_{u_{k}}$. In particular, we include $(r_{k},l_{k},y_{k-1},a_{k-1},l_{k-1},r_{k-1})$ as well as baseline covariates. Note, $u_{k-1}$ is a subset of $l_{k-1}$ and is observed for $k=3$ and must first be estimated for $k>3$. Since a single estimate corresponds to just one realization from the distribution of $U_{k}$, we simulate multiple draws (say 5) of $U_{k}$ for $k>2$, and an overall estimate of the intervention effect is obtained by averaging over the single estimates.
Positivity violations: To assess robustness of results from positivity violations (C3) we discarding subjects with insufficient overlap. To identify these subjects we implemented a generalized propensity score approach, where the propensity score models at times $t=2,\ldots,T$ were fitted using L-DART. Subjects whose sBP was above the given threshold were excluded if their generalized propensity score were out of the range of the generalized propensity score among subjects with sBP within optimal levels.
Missingness for the outcome: Previous studies indicate that the dropout may be MNAR for the outcome [16]. As such, we introduce a sensitivity parameter $\gamma_{t}$ to identify the distribution for dropouts among survivors, to allow for deviations from the MARS assumption (C4). We assume $p(y_{t}|\bar{a}_{t},\bar{l}_{t},r_{t}=0,\bar{s}_{t}=1,\bar{y}_{t-1})=p(y_{t}-\gamma_{t}|\bar{a}_{t},\bar{l}_{t},\bar{r}_{t}=1,\bar{s}_{t}=1,\bar{y}_{t-1})$ for all $t>0$. Our prior belief is that $\gamma_{t}<0$, reflecting a negative shift in memory after a subject drops out. Here, the prior is specified as $\gamma_{t}\sim Triangular(L_{\gamma_{t}},0,L_{\gamma_{t}}),$ where we assume the lower bound not to be bigger than the estimated difference in memory change from baseline to follow-up between dropouts and completers, when adjusting for age and sex. That is, $E[EMS_{i2}-EMS_{i1}|age_{i1},sex_{i1},r_{i3}=1]-E[EMS_{i2}-EMS_{i1}|age_{i1},sex_{i1},r_{i3}=0]$, where $EMS_{it}$ is the episodic memory score at time $t$.
Results: To ease presentation we divided the sample into two age groups; middle-aged, i.e. $age_{iT}<65$, and a older group, i.e $age_{iT}\geq 65$. First, the results were consistent with the results obtained using default assumptions for both age groups when taking into account a unmeasured time-varying confounder, i.e. cholesterol levels. We further investigated the impact of possible positivity violations on the results. For this analysis, subjects with limited overlap ($N=61$) was first discarded. The results did not differ significantly from the main analysis either. Last, the posterior mean episodic memory score was attenuated both under the natural course and under the intervention when using the MNARS assumption instead of the MARS on the outcome. However, the estimates of the $PCIE$ did not change. See detailed results in Table A4.
8 Conclusions
Using data from a prospective cohort study on aging, cognition and dementia, we implemented a hypothetical incremental threshold intervention to investigate whether blood pressure monitoring at more optimal levels over 15 years was associated with memory improvements. In our study, there were no significant prognostic- or etiological effects across all ages. These findings provide important information about long-term blood pressure control in cognitive aging, indicating that sBP interventions likely do not have a strong effect on memory neither at the population level nor at the individual level. However, more research from long-term clinical trials on whether lowering sBP can prevent cognitive decline is needed to draw firm conclusions.
[…]
We propose a Bayesian semi-parametric approach for the proposed incremental intervention, extending prior research in several ways. First, our approach introduces an incremental intervention for prospective cohort data with attrition, offering a viable alternative when clinical trials are not available or feasible to conduct. second, we offer two types of mortal cohort inferences, providing both prognostic and etiological effect estimates, both of which are critical for public health practice, especially in older populations. Last, our modeling approach employs soft BART models with a novel sparsity-inducing hyperprior for longitudinal data, promoting parsimony in longitudinal prediction settings and demonstrating good performance in simulations.
Acknowledgments
This work is funded by Forte Dnr 2019-01064 to MJ, and NIH RO1 HL166324 and HL158963 to MJD. This research was conducted using the resources of High Performance Computing Center North (HPC2N). The Betula Project is supported by Knut and Alice Wallenberg foundation (KAW) and the Swedish Research Council (K2010-61X-21446-01).
Supplementary material
The Supplementary Appendices A and B provide details for the extended G-formula and identification of the SAIE. Computational details for the longitudinal Dirichlet prior and the G-computation algorithm are given in Appendix C. Supplementary simulation results are given in Appendix D and supplementary Betula results are given in Appendix E.
Supplementary material - Long-term cognitive effects of an incremental blood pressure intervention in a mortal cohort
Josefsson, Maria, Karalija, Nina, & Daniels, Michael J.
A. The Extended G-formula
Below we provide details of the Extended G-formula as introduced in Section 4.2.
$\int_{\bar{s}_{t}}\int_{\bar{r}_{t}}\int_{\bar{l}_{t}}\int_{\bar{a}_{t}^{*}}\int_{\bar{a}_{t}}\int_{\bar{y}_{t-1}}E[Y_{T}|\bar{y}_{T-1},\bar{a}_{T},\bar{l}_{T},\bar{r}_{T},\bar{s}_{T}=1]$
×$\quad\times\prod_{k=1}^{T}p^{int^{*}}(a_{k}|\bar{y}_{k-1},a^{*}_{k},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)$
×$\quad\times p(a^{*}_{k}|\bar{y}_{k-1},d_{k-1},\bar{l}_{k},\bar{s}_{k}=1)$
×$\quad\times p(l_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k},\bar{s}_{k}=1)$
×$\quad\times p(r_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k}=1)$
×$\quad\times p(s_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1)$
×$\quad\times p(y_{k-1}|\bar{y}_{k-2},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1).$ (A.1)
B. Identification of the SAIE under a hypothetical intervention based on the natural value of treatment
Identification of the SAIE under a hypothetical intervention based on the NVT under a stochastic Monotonicity assumption is presented below. First, for the two contrasting regimes $g^{nc}$ and $g^{int^{*}}$, we have that the expected potential outcome for the principal strata becomes,
$E[Y_{T}(g^{int^{*}})-Y_{T}(g^{nc})\mid\bar{S}(g^{nc})=S_{T}(g^{int^{*}})=1]=$
$\quad PCIE+\Delta\left\{\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)\right\}\left(1-U^{-1}\right),$ (A.2)
where $U=\min\left\{1,\frac{\psi^{g^{int^{*}}}}{\psi^{g^{nc}}}\right\}$. Identification of [S0.Ex14] is obtained by using the law of total probability (ltp) and some algebra. First, for the intervention regime $g^{int^{*}}$, and similarly for $g^{nc}$, we have that
$E[$ $Y_{T}(g^{int^{*}}))\mid S_{T}(g^{int^{*}})=1]=$
$\quad E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{nc})=S_{T}(g^{int^{*}})=1]+\Pr[S_{T}(g^{nc})\neq 1\mid S_{T}(g^{int^{*}})=1]\times$
$\quad\left\{E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{nc})=1,S_{T}(g^{int^{*}})\neq 1]-E[Y_{T}(g^{int^{*}})\mid\bar{S}(g^{nc})=S_{T}(g^{int^{*}})=1]\right\}.$
Using A6 and solving the above equation for $E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{nc})=S_{T}(g^{int^{*}})=1]$ we have that
$E[$ $Y_{T}(g^{int^{*}})\mid\bar{S}(g^{nc})=S_{T}(g^{int^{*}})=1]=$
$E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{int^{*}})=1]+\Delta\times\left\{1-\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]\right\},$
where $\Delta$ is a sensitivity parameter when comparing the difference in potential outcomes when comparing the principal strata to the strata where subjects were to live under the hypothetical intervention but not under the natural course. Hence,
$E[Y_{T}(g^{int^{*}})-Y_{T}(g^{nc})\mid\bar{S}(g^{nc})=S_{T}(g^{int^{*}})=1]$
$\quad=E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{int^{*}})=1]+\Delta\times\left\{1-\Pr[S_{T}(g^{nc})=1\mid S_{T}(g^{int^{*}})=1]\right\}$
$\qquad-E[Y_{T}(g^{nc})\mid S_{T}(g^{nc})=1]-\Delta\times\left\{1-\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]\right\}$
$\quad=E[Y_{T}(g^{int^{*}})\mid S_{T}(g^{int^{*}})=1]-E[Y_{T}(g^{nc})\mid S_{T}(g^{nc})=1]+$
$\qquad\Delta\times\left\{\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]-\Pr[S_{T}(g^{nc})=1\mid S_{T}(g^{int^{*}})=1]\right\}$
$\quad=PCIE+\Delta\times\left\{\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]-\Pr[S_{T}(g^{nc})=1\mid S_{T}(g^{int^{*}})=1]\right\}$
Moreover, by using A5 (Stochastic monotonicity) and Bayes theorem we have that
$\Pr[S_{T}(g^{int^{*}})=1\mid S_{T}(g^{nc})=1]=\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)$ and $\Pr[S_{T}(g^{nc})=1\mid S_{T}(g^{int^{*}})=1]=\frac{\psi^{g^{nc}}}{\psi^{g^{int^{*}}}}\left[\psi^{g^{int^{*}}}+\lambda\left(U-\psi^{g^{int^{*}}}\right)\right]$, where $\lambda$ is a sensitivity parameter, and $0\leq\lambda\leq 1$. Hence, the SAIE simplifies to (A.2).
For $g\in(g^{int^{*}},g^{nc})$ and using A1-A4, we have that
$E[Y_{T}(g)\mid S_{T}(g)=1]$ $=\frac{E[Y(g),S_{T}(g)=1]}{\Pr[S_{T}(g)=1]}$
$=\frac{E_{H^{g}}[E(Y_{T},S_{T}=1\mid H^{g})]}{E_{H^{g}}[\Pr(S_{T}=1\mid H^{g})]}.$ (A.3)
[…]
$E$ $[Y_{T}(g^{nc}),S_{T}(g^{nc})=1]=E_{H^{g^{nc}}}\{E[Y_{T}(g^{nc}),S_{T}(g^{nc})=1\mid H^{g^{nc}}]\}$
$=\int\ldots\int E[Y_{T}(g^{nc})\mid\bar{S}_{T}=1,\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{L}_{T}=\bar{l}_{T},\bar{R}_{T}=\bar{r}_{T}]$
×$\quad\times\Pr[\bar{S}_{T}(g^{nc})=1\mid\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{L}_{T-1}=\bar{l}_{T-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(Y_{k-1}(g^{nc})=y_{k-1}\mid\bar{Y}_{k-2}=\bar{y}_{k-2},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p(A_{k}(g^{nc})=a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(L_{k}(g^{nc})=l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(R_{k}(g^{nc})=r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
Exchangeability
$=\int\ldots\int E[Y_{j}(g^{nc})\mid\bar{S}_{t}=1,\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t}=z_{t},\bar{L}_{j}=\bar{l}_{j},\bar{R}_{T}=\bar{r}_{T}]$
×$\quad\times\Pr[\bar{S}_{t}(g^{nc})=1\mid\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=z_{t-1},\bar{L}_{t-1}=\bar{l}_{t-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(y_{k-1}\mid\bar{Y}_{k-2}=\bar{y}_{k-2},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p(a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
Consistency
$=\int\ldots\int E[Y_{j}\mid\bar{S}_{t}=1,\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t}=z_{t},\bar{L}_{j}=\bar{l}_{j},\bar{R}_{T}=\bar{r}_{T}]$
×$\quad\times\Pr[\bar{S}_{t}=1\mid\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=z_{t-1},\bar{L}_{t-1}=\bar{l}_{t-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(y_{k-1}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p(a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
$=E_{H^{g^{nc}}}\{E[Y,S_{T}=1\mid H^{g^{nc}}]\}=E[Y,S_{T}=1].$
A similar argument shows that $\Pr[S_{T}(g^{nc})=1]=E_{H_{int}}\{\Pr[S_{T}(g^{nc})=1\mid H_{int}]\}=E_{H_{int}}\{\Pr[S_{T}=1\mid H_{int}]\}=\Pr[S_{T}=1]$, and therefore,
$\frac{E[Y(g^{nc}),S_{T}(g^{nc})=1]}{\Pr[S_{T}(g^{nc})=1]}=E_{g^{nc}}[Y\mid S_{T}=1].$ (A.4)
Hence, it follows that (A.4) can be estimated using the g-formula conditioning on survival at a specific time point as described in (S0.Ex2) in the main text.
For regime $g^{int^{*}}$, $E_{H^{g^{int^{*}}}}[E(Y_{T},S_{T}=1\mid H^{g^{int^{*}}})]$ is also obtained by marginalizing over the distributions of the set of temporally preceding variables $H^{g^{int^{*}}}$. However, the extended g-formula must be used for a intervention that depends on the NVT [31]. As such, for regime $g^{int^{*}}$ we have that
$E$ $(Y_{T}(g^{int^{*}}),S_{T}(g^{int^{*}})=1)=E_{H^{g^{int^{*}}}}[E(Y_{T}(g^{int^{*}}),S_{T}(g^{int^{*}})=1)\mid H^{g^{int^{*}}}]$
$=\int\ldots\int E[Y_{T}(g^{int^{*}})\mid\bar{S}_{T}=1,\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{L}_{T}=\bar{l}_{T},\bar{R}_{T}=\bar{r}_{T}]$
×$\quad\times\Pr[\bar{S}_{T}(g^{int^{*}})=1\mid\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{L}_{T-1}=\bar{l}_{T-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(Y_{k-1}(g^{int^{*}})=y_{k-1}\mid\bar{Y}_{k-2}=\bar{y}_{k-2},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p^{int^{*}}(a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},A^{*}_{k}=a^{*}_{k},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(A^{*}_{k}(g^{int^{*}})=a^{*}_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(L_{k}(g^{int^{*}})=l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(R_{k}(g^{int^{*}})=r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}^{*}_{T}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
=$\ldots$= …
$\ldots$
Exchangeability
$=\int\ldots\int E[Y_{T}(g^{int^{*}})\mid\bar{S}_{T}=1,\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{A}_{T}=z_{T},\bar{L}_{T}=\bar{l}_{T}]$
×$\quad\times\Pr[\bar{S}_{T}(g^{int^{*}})=1\mid\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{A}_{T-1}=a_{T-1},\bar{L}_{T-1}=\bar{l}_{T-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(y_{k-1}\mid\bar{Y}_{k-2}=\bar{y}_{k-2},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p^{int^{*}}(a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},A^{*}_{k}=a^{*}_{k},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(a^{*}_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}^{*}_{T}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
Consistency
$=\int\ldots\int E[Y_{T}\mid\bar{S}_{T}=1,\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{A}_{T}=a_{T},\bar{L}_{T}=\bar{l}_{T},\bar{R}_{T}=\bar{r}_{T}]$
×$\quad\times\Pr[\bar{S}_{T}=1\mid\bar{Y}_{T-1}=\bar{y}_{T-1},\bar{A}_{T-1}=a_{T-1},\bar{L}_{T-1}=\bar{l}_{T-1},\bar{R}_{T-1}=\bar{r}_{T-1}]$
×$\quad\times\prod_{k=1}^{T}p(y_{k-1}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k-1}=1)$
×$\qquad\times p^{int^{*}}(a_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},A^{*}_{k}=a^{*}_{k},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(a^{*}_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k}=\bar{l}_{k},\bar{R}_{k}=\bar{r}_{k},\bar{S}_{k}=1)$
×$\qquad\times p(l_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=z_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)$
×$\qquad\times p(r_{k}\mid\bar{Y}_{k-1}=\bar{y}_{k-1},\bar{A}_{k-1}=\bar{a}_{k-1},\bar{L}_{k-1}=\bar{l}_{k-1},\bar{R}_{k-1}=\bar{r}_{k-1},\bar{S}_{k}=1)d\bar{y}_{T-1}d\bar{a}^{*}_{T}d\bar{a}_{T}d\bar{l}_{T}d\bar{r}_{T}$
$=E_{H^{g^{int^{*}}}}[E(Y_{T},S_{T}=1\mid H^{g^{int^{*}}})]=E(Y_{T},S_{T}=1).$
Similarly as for $g^{nc}$ we have that $\Pr[S_{T}(g^{int^{*}})=1]=E_{H^{g^{int^{*}}}}\{\Pr[S_{T}(g^{int^{*}})=1\mid H^{g^{int^{*}}}]\}=E_{H^{g^{int^{*}}}}\{\Pr[S_{T}=1\mid H^{g^{int^{*}}}]\}=\Pr[S_{T}=1]$, and therefore,
$\frac{E[Y(g^{int^{*}}),S_{T}(g^{int^{*}})=1]}{\Pr[S_{T}(g^{int^{*}})=1]}=E_{g^{int^{*}}}[Y\mid S_{T}=1].$ (A.5)
Hence, it follows that (A.5) can be estimated using the extended g-formula conditioning on survival at a specific time point as described in (S0.Ex7) in the main text. Moreover, by combining the results above we show that the SAIE become (A.2).
C. Computational details
Posterior computations
[…]
G-computation algorithm
  • 1.
    Modeling the observed data: For $t=0,\ldots,T$, sample from the observed data posteriors for the parameters of the conditional distributions of $y_{t}$, $a_{t}$, $l_{t}$, $r_{t}$ and $s_{t}$ using L-DART. Note, we implement the Bayesian bootstrap [5] to integrate over the distribution of baseline confounders without missingness.
  • 2.
    Pseudo data: For each posterior sample of the parameters of the L-DART models in Step 1, sequentially sample pseudo data, $D^{nc}$ and $D^{int^{*}}$, of size $N^{p}$ for the two contrasting regimes. Under the hypothetical intervention, to obtain pseudo data for $a_{t}$, pseudo NVT data is first sampled from the observed posterior distribution $a^{*}_{t}\sim N\left(\mu_{a_{t}},\sigma_{t}^{2}\right)$ to compute $a_{t}=a^{*}_{t}+\delta_{t}I\{a^{*int}_{t}\geq\tau\}$, where a set of the shift parameter $\delta_{t}$ is first sampled from the associated prior distribution.
  • 3.
    G-computation: Implement G-computation, as described in Section 3 for the two contrasting regimes using the pseudo data, $D^{nv}$ and $D^{int*}$, and sensitivity parameters, $\bar{\delta}_{t}$ from Step 2 to obtain posterior estimates, $\hat{\mu}^{int^{*}}$ and $\hat{\mu}^{nc}$, of $E_{H_{int^{*}}}\{E[Y_{T}|S_{T}=1,H_{int^{*}}]\}$ and $E_{H_{nc}}\{E[Y_{T}|S_{T}=1,H_{nc}]\}$ respectively. For the SAIE this also involves implementing G-computation for computing predicted survival at time $T$, i.e. $\hat{\psi}^{g}=\prod_{t=2}^{T}\hat{s}^{g}_{t}$ for the two contrasting regimes $g\in(g^{int^{*}},g^{nc})$.
  • 4.
    Computation of the PCIE and the SAIE: Implement Monte Carlo integration using the pseudo data for the two regimes to compute one posterior sample of the
    $PCIE=\frac{\sum_{i}\hat{\mu}^{int^{*}}_{iT}s^{int^{*}}_{iT}}{\sum_{t}^{T}\sum_{i}s^{int^{*}}_{iT}}-\frac{\sum_{i}\hat{\mu}^{nc}_{iT}s^{nc}_{iT}}{\sum_{t}^{T}\sum_{i}s^{nc}_{iT}}$
    and the
    $SAIE=PCIE+\Delta\left\{\hat{\psi}^{g^{int^{*}}}+\lambda\left(U-\hat{\psi}^{g^{int^{*}}}\right)\right\}\left(1-U^{-1}\right),$
    where $U=\min\left\{1,\frac{\hat{\psi}^{g^{int^{*}}}}{\hat{\psi}^{g^{nc}}}\right\}$, and one must also sample one set of the sensitivity parameters $\Delta$ and $\lambda$.
  • 5.
    Repeat step 2 - 4 for each posterior sample.
D. Simulation results
Table A1: Results in terms of % bias from the simulation study under the different scenarios; A-E using either 1/10/100 regression trees. For scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, C: $\bar{c}=(1,0,0,0)$, D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, and E: $\bar{c}=(0.25,0.25,0.25,0.25)$. We compare five different BART models; L-DART: our approach, H-BART: the standard BART algorithm, H-DART: BART with the sparsity-inducing Dirichlet prior, S-DART: the standard soft BART algorithm, and S-DART-lag1: SoftBart incorporating only the predictors from the last time point.
A B C D E
n=1500
L-DART-1 -0.0 -0.1 -0.0 -0.1 -0.2
S-DART-1 0.0 -0.1 -0.1 -0.1 -0.2
S-DART_l1-1 -0.1 -0.1 -0.0 -0.0 -0.1
H-DART-1 -0.1 -0.1 -0.0 0.0 -0.1
H-BART-1 0.0 -0.0 -0.1 0.0 -0.1
L-DART-10 -0.1 -0.0 0.2 -0.0 -0.1
S-DART-10 -0.1 0.0 0.2 -0.0 -0.1
S-DART_l1-10 -0.0 -0.1 0.2 0.0 -0.1
H-DART-10 -0.1 0.0 0.2 -0.0 -0.0
H-BART-10 -0.1 -0.0 0.2 0.0 -0.1
L-DART-100 -0.1 -0.2 -0.0 -0.1 0.2
S-DART-100 -0.1 -0.2 -0.0 -0.0 0.2
S-DART_l1-100 -0.0 -0.2 -0.0 -0.0 0.1
H-DART-100 -0.0 -0.1 0.0 -0.0 0.2
H-BART-100 -0.0 -0.1 0.1 -0.0 0.2
n=500
L-DART-1 -0.2 -0.2 0.1 -0.2 0.0
S-DART-1 -0.2 -0.1 0.1 -0.2 0.0
S-DART_l1-1 0.0 -0.2 0.1 -0.2 0.0
H-DART-1 -0.0 -0.0 0.3 0.1 0.1
H-BART-1 0.0 0.1 0.2 0.0 0.2
L-DART-10 -0.1 -0.1 -0.0 0.1 -0.1
S-DART-10 -0.1 -0.2 0.0 0.0 -0.1
S-DART_l1-10 -0.1 -0.0 0.1 0.1 0.0
H-DART-10 0.1 -0.0 0.2 0.2 -0.1
H-BART-10 0.1 0.1 0.3 0.1 -0.0
L-DART-100 -0.0 -0.1 -0.3 -0.1 0.1
S-DART-100 -0.0 0.0 -0.3 -0.1 0.1
S-DART_l1-100 0.0 -0.1 -0.2 -0.1 0.1
H-DART-100 0.0 0.1 -0.1 -0.0 0.1
H-BART-100 0.0 0.1 -0.1 -0.0 0.2
n=100
L-DART-1 -0.3 0.1 -0.9 -0.8 -0.5
S-DART-1 -0.3 0.0 -0.6 -0.7 -0.5
S-DART_l1-1 0.0 0.1 -0.7 -0.4 -0.6
H-DART-1 0.2 0.6 -0.0 -0.1 -0.4
H-BART-1 0.5 0.6 -0.0 -0.3 -0.4
L-DART-10 -0.5 -0.2 -0.9 -0.5 -1.1
S-DART-10 -0.5 -0.1 -0.5 -0.5 -1.3
S-DART_l1-10 -0.4 -0.1 -0.8 -0.2 -0.8
H-DART-10 0.2 0.4 -0.5 0.1 -1.0
H-BART-10 0.2 0.5 -0.8 -0.0 -1.0
L-DART-100 -0.6 -0.7 -0.2 -0.4 -0.6
S-DART-100 -0.6 -0.6 0.1 -0.4 -0.6
S-DART_l1-100 -0.1 -0.5 -0.3 -0.3 -0.3
H-DART-100 0.1 -0.2 0.2 0.0 -0.2
H-BART-100 0.0 -0.3 0.2 0.1 -0.1
Table A2: Results from the simulation study under the different scenarios; A - E, of BART models fitted using 1 regression tree. For scenario A: $\bar{c}=(1,0.75,0.5,0.25)$, B: $\bar{c}=(1,0.75,0.5,0.25)^{3}$, C: $\bar{c}=(1,0,0,0)$, D: $\bar{c}=(1,0.75,0.5,0.25,0,0,0,0)$, E: $\bar{c}=(0.25,0.25,0.25,0.25)$. The results are presented in terms of RMSE and 95% coverage probabilities (CP). We compare five different BART models; L-DART: our approach, H-BART: the standard BART algorithm, H-DART: BART with the sparsity-inducing Dirichlet prior, S-DART: the standard soft BART algorithm, and S-DART-lag1: SoftBart incorporating only the predictors from the last time point.
A B C D E
RMSE CP RMSE CP RMSE CP RMSE CP RMSE CP
n=1500
L-DART-1 2.67 0.95 2.05 0.95 1.52 0.95 2.77 0.95 1.68 0.95
S-DART-1 2.69 0.95 2.09 0.95 1.60 0.95 2.79 0.95 1.69 0.95
S-DART_l1-1 5.43 0.86 2.69 0.93 1.50 0.95 5.43 0.86 2.69 0.75
H-DART-1 6.90 0.78 4.78 0.80 3.64 0.81 6.92 0.77 2.89 0.70
H-BART-1 6.98 0.77 4.88 0.78 3.73 0.80 7.01 0.74 2.89 0.72
n=500
L-DART-1 3.59 0.95 2.78 0.95 2.13 0.95 3.80 0.94 2.11 0.95
S-DART-1 3.61 0.95 2.86 0.95 2.45 0.95 3.86 0.95 2.15 0.95
S-DART_l1-1 5.67 0.86 3.11 0.92 2.12 0.95 5.68 0.86 2.88 0.71
H-DART-1 7.29 0.75 5.15 0.76 4.08 0.75 7.33 0.72 2.98 0.73
H-BART-1 7.31 0.75 5.16 0.76 4.11 0.74 7.38 0.71 2.98 0.74
n=100
L-DART-1 5.37 0.95 4.25 0.94 3.65 0.94 5.82 0.95 2.93 0.91
S-DART-1 5.42 0.95 4.63 0.93 4.23 0.90 6.07 0.95 2.94 0.91
S-DART_l1-1 6.58 0.85 4.36 0.90 3.57 0.92 6.58 0.85 3.18 0.76
H-DART-1 7.74 0.75 5.69 0.78 4.65 0.81 7.87 0.72 3.13 0.91
H-BART-1 7.72 0.76 5.68 0.79 4.61 0.81 7.88 0.72 3.15 0.91
E. Betula results
Table A3: Baseline demographics for the middle-aged and old group in the Betula study, presented as mean (standard deviation) for the continuous variables and % (N) for the dichotomous variables.
Middle aged Older
Age 42.5 (5.6) 62.5 (5.6)
Male 0.46 (184) 0.50 (198)
Education 12.6 (3.8) 8.5 (3.2)
sBP 124.8 (15.6) 147.7 (22.0)
dBP 79.6 (10.0) 85.6 (9.3)
Smoker 0.57 (230) 0.53 (212)
BMI 24.3 (3.3) 25.7 (3.5)
Cholesterol 52.0 (19.2) 60.8 (19.3)
Diabetes type 2 0.01 (6) 0.07 (28)
Episodic memory (correct) 0.53 (0.12) 0.42 (0.11)
Table A4: Results for the $PCIE$ from the sensitivity analyses of the Betula data using the proposed approach. MARS refers to the default settings as described in Section 7. Unmeasured confounding refers to a sensitivity analysis for assumption A2, Positivity refers to a sensitivity analysis for A3, and MNARS refers to a sensitivity analysis A4b. See Section 7.2 for further details.
Middle aged Older
MARS 0.4 (-0.6; 1.4) 1.3 (-0.6; 3.3)
Sensitivity analysis
Unmeasured confounding 0.3 (-0.6; 1.35) 1.3 (-0.6; 3.3)
Positivity 0.3 (-0.6; 1.24) 1.1 (-0.7; 3.0)
MNARS 0.4 (-0.6; 1.38) 1.3 (-0.7; 3.3)